In air an object weight 15N. When immersed in water it weight 12N. When immersed in another liquid it weight 13N. Calculate the density of the object and that of the other liquid.


1
Expert's answer
2021-06-21T16:27:34-0400

Upthrust acting on the block in water

=15N−12N=3N = ( volume of solid immersed )( density of liquid ) g = V (1) 10

where V is the volume of the block , so V= "\frac{3}{10}=0.3cm^3"


Block weighs 15N in air implies 15=mg implies m= "\frac{15}{10}=1.5g"

now density = "\frac{m}{v}=\frac{15}{0.3}=5g" /cm3

so density of block = 5 gm/cm3

hence , relative density of the block =

"\frac{density of block }{ density of water}=\frac{5}{1}=5gm/cm^3"


now Upthrust acting on the block in the liquid =15N−13N=2N= (volume of solid immersed ) ( density of liquid ) g=Vρg=0.3( ρ)g ,

implies ρ of the liquid = "\frac{2}{3}gm/cm^3"


so relative density of the liquid ="\frac{density of liquid }{ density of water}=\frac{\frac{2}{3}}{1}=\frac{2}{3}"